OpenStudy (anonymous):

Josue's overtime pay is $6 an hour more than his regular pay. He worked 7 hours at his regular wage and 5 hours at his overtime wage. He earned $138. What is Josue's regular wage per hour? $6 per hour $9 per hour $12 per hour $15 per hour

OpenStudy (anonymous):

HELPP PLEASE

OpenStudy (ganpat):

let regular pay be x So overtime is = x +6 so 7x + 5 (x + 6) = 138 can u solve further ??

OpenStudy (anonymous):

YES

OpenStudy (ganpat):

7x + 5 (x + 6) = 138 7x + 5x + 30 = 138 12x = 138 - 30 12x = 108 x = 9 $ = regular pay and overtime = 9 + 6 = 15 $... Nice too see ur confidence :)

OpenStudy (anonymous):

WOW